# CODA: Repurposing Continuous VAEs for Discrete Tokenization

### v1

@@ 5.3 Ablation Studies
Results in Tab. 4(b) demonstrate that our approach enables stable scaling of the codebook size from $1024$ to $65536$, leading to a consistent improvement in tokenizer performance, with rFID steadily decreasing from $2.07$ to $1.34$. This highlights the effectiveness of our method in fully utilizing an expanded codebook to refine discrete representations without encountering common issues such as codebook collapse or underutilization.
@@
We argue that normalization used in the attention quantization modules play a positive role in better assignment, leading to improved reconstruction results. As results in Tab. 4(c) demonstrate, projecting the query-key pairs into a normed space improves rFID from $1.87$ to $1.34$. Other types of normalization, ex. LayerNorm is also capable of achieving similar results compared to using RMSNorm.

### v2

@@ 5.3 Ablation Studies
Results in Tab. 4(b) demonstrate that our approach enables stable scaling of the codebook size from $1024$ to $65536$, leading to a consistent improvement in tokenizer performance, with rFID steadily decreasing from $2.07$ to $1.34$. This highlights the effectiveness of our method in fully utilizing an expanded codebook to refine discrete representations without encountering common issues such as codebook collapse or under-utilization.
@@
We argue that normalization used in the attention quantization modules play a positive role in better assignment, leading to improved reconstruction results. As results in Tab. 4(c) demonstrate, projecting the query-key pairs into a normed space improves rFID from $1.87$ to $1.34$. Other types of normalization, e.g. LayerNorm is also capable of achieving similar results compared to using RMSNorm.
@@ 6 Conclusion
Acknowledgement
This work is supported in part by the National Key R&D Program of China (2024YFB4708200), the National Natural Science Foundation of China (U24B20173, 62321005), and the Scientific Research Innovation Capability Support Project for Young Faculty (ZYGXQNJSKYCXNLZCXM-I20).